# Deep learning interpretability analysis for carbon star identification in DR3

### v1

@@ 6 Conclusions
In this work, we have demonstrated the capability of the proposed βGaiaNetβ model in classifying carbon stars within the CSTAR_XPM sample. Additionally, we have provided compelling feature attribution explanations for the modelβs outputs using the SHAP method. This βmagicβ empowers the originally incomprehensible black-box model with the ability to proactively βspeak upβ and make its discriminations traceable. Our main results and conclusions are as follows:
@@
The method presented in C2023 for calculating the strength of molecular bands has proven effective in selecting carbon stars with strong CN features and temperatures below 3,800K. However, it lacks effectiveness in identifying warmer carbon stars with weaker CN features and may introduce contamination. In contrast, the proposed βGaiaNetβ model demonstrates superior capability in distinguishing between carbon and non-carbon stars with good robustness from the training datasets, particularly in identifying carbon stars with weak CN features from CSTAR_XPM_nonG. Compared to four conventional machine learning methods, the βGaiaNetβ model exhibits an average accuracy improvement of approximately 0.3% on the validation set, with the highest accuracy reaching up to 100%. The flexible parameter architecture provides the model with enhanced fitting capability and stability achieved through the incorporation of operations such as 1Γ1 Convolution, Global Average Pooling, Batch Normalization (BN), and Dropout layers. Furthermore, Optimised training methods including momentum and weight decay were employed. By utilizing deep learning models with varying convolutional kernels during the 1D convolution process, the βGaiaNetβ model effectively captures both local and global features from low-resolution smoothed spectra. Coupled with the nonlinear capabilities provided by an appropriate modelβs depth, the model possesses natural advantages in handling βs XP spectra.
@@
We utilized SHAP for feature attribution interpretation on the βGaiaNetβ classification model based on positive sample. We provided clear feature heatmaps for each spectrum, which serve as a basis for judgment and identification of each spectrum. Based on the statistical analysis of SHAP values from the spectra, we found that the C features are almost non-distinguishing for the identification of cold carbon stars due to their weak signal at the blue end of low flux values and the minimal presence in many cold carbon stars. We summarised five key spectral feature regions: 6780-7460βΓ , 7600-7820βΓ , 7900-8360βΓ , 8820-9060βΓ , and 9140-9560βΓ . Among them, the 7900-8360βΓ Β and 9140-9560βΓ Β regions are the most obvious trough areas. These two regions contribute over 80% due to dense CN molecular absorption. The 7600-7820βΓ Β and 8820-9060βΓ Β regions contribute about 16%, which are are also important distinguishing features due to the peak areas from the strong CN molecular absorption on both sides. The 6780-7460βΓ Β region contains relatively weak peaks and troughs, making the smallest contribution less than 4. We hypothesize that this is related to the spectral values, which seems to be addressable through improved spectral normalization methods. These results suggest that our model can effectively learn and capture the critical distinguishing features that align well with CN molecular absorption lines and visual identifications by astronomers when identifying carbon star spectra. These five key feature regions will be essential for future systematic mining of carbon stars from βs XP spectra.
There are 451 new carbon star candidates finally found from CSTAR_XPM_nonG using the trained classification model. Compared with the method of simply calculating band head strength, our approach effectively identifies more carbon stars with weaker CN features. Compared to the golden sample carbon stars, these new candidates exhibit relatively weaker CN molecular band features and occupy the positions of weak CN feature carbon stars illustrated in Fig. 21 presented in C2023. This means that the carbon abundances in their atmospheres may be relatively low, with C/O ratios not significantly greater than 1. These candidates are mainly concentrated in the Galactic inner disk regions near the center. Additionally, we also noted that most of them have higher temperatures (smaller ) and few of them are labelled as LPVs by Lebzelter etΒ al. (2023), indicating the possibility that they could be mostly extrinsic and less evolved carbon stars. This is similar to the findings in Sanders & Matsunaga (2023). Moreover, our candidates exhibit stronger negative contributions caused by CaβII absorption features near 8500βΓ Β compared to the positive sample from the statistical analysis of SHAP values. Therefore, some of these stars are also likely to be AGB stars with relatively mild third dredge-up processes and higher metallicity, while also not excluding the possibility that they were contaminated by a carbon rich environment near the Galactic center (Abia etΒ al. 2002).
@@
We cross-matched the 451 new candidates with LAMOST DR10, resulting in a total of 4 common sources. Among them, 2 were identified as carbon stars by the LAMOST 1D pipeline. By cross-matching with the SIMBAD database, we obtained 73 common sources. Among these, 28 had already been reported as carbon stars or carbon star candidates in previous work. A reference sample of 1,289 LPVs was constructed to compare the results in Sanders & Matsunaga (2023). In this work, we have discovered 40 potentially carbon stars from it, and they identified 26 C-rich stars including only 6 stars that are not in our candidates. This is due to the significantly different spectral features of the six stars from the positive sample, which results in their confidence lower than 0.5. These results indicate that the reliability and completeness of our candidates are high. Furthermore, the assigned confidence and SHAP values for each spectrum were used to quantify the modelβs output, revealing a strong correlation with the band head strengths of R773.3 and R895.0. This suggests that the classification model primarily focuses on the crucial CN features, and the learned features make sense. Hence, our model can efficiently learn the key features between carbon and non-carbon stars in the CSTAR_XPM sample, and the modelβs reliability is further validated.
@@
Compared to the positive sample, Most CSTAR_XPM_nonG exhibit distinctly different distributions (much weaker) on the band head strength diagrams. However, our trained model still identifies 451 credible potential carbon stars within it, which also demonstrates the generalization ability of the model. These carbon stars exhibit higher temperatures and weaker CN features, making them undetectable by the C2023 method. Due to the limited diversity of the training positive sample, there are still some carbon stars with weaker CN features that are not recognized by our algorithm. Their probabilities given by the modelβs output do not exceed the threshold of 0.5 but significantly larger than 0, thus the proportion of candidates in CSTAR_XPM_nonG could be higher, especially the G and K type stars, predictably if we set a lower threshold. we roughly estimate that the loss percentage is at least about 5%. In the next work, we will make reasonable improvements to the classification algorithm and incorporate more carbon stars with weaker CN bands into positive sample to search for and identify more potential warm candidates from CSTAR. We will then expand the algorithm to the entire XP spectral dataset, thereby significantly enriching the number of known carbon stars. Indeed it would be interesting to perform further classification and meaningful analysis. Given the resolution of βs XP spectra, our method can be directly transferred to the upcoming China Space Station Telescope (CSST), especially for the interpretation of different types of stars whose spectral properties are still unclear.
@@ Acknowledgments
Appendix A The Heat Map of Six diffent stars
@@ Appendix B New Carbon star candidates from CSTAR_XPM_nonG
The catalog has 451 entries, which are all the most likely carbon star candidates selected by our model. The model obtains this result by training positive sample selected from CSTAR_XPM_G and negative sample selected from CSTAR_XPM_non_G. All candidates are sorted in descending order of confidence, and we show the top 100 candidates.

### v2

@@ 6 Conclusions
In this work, we have demonstrated the capability of the proposed βGaiaNetβ model in classifying carbon stars within the CSTAR_XPM sample. Additionally, we have provided compelling feature attributional interpretations for the modelβs outputs using the SHAP method. This βmagicβ empowers the originally incomprehensible black-box model with the ability to proactively βspeak upβ and make its discriminations traceable. Our main results and conclusions are as follows:
@@
The method presented in C2023 for calculating the strength of molecular bands has proven effective in selecting carbon stars with strong CN features. However, it lacks effectiveness in identifying hotter carbon stars with weaker CN features and may introduce contamination. In contrast, the proposed βGaiaNetβ model demonstrates superior capability in distinguishing between carbon and non-carbon stars robustly from the training datasets, particularly in identifying carbon stars with weak CN features from CSTAR_XPM_nonG. Compared to four conventional machine learning methods, the βGaiaNetβ model exhibits an average accuracy improvement of approximately 0.3% on the validation set, with the highest accuracy reaching up to 100%. The flexible parameter architecture provides the model with enhanced fitting capability and stability achieved through the incorporation of operations such as 1Γ1 Convolution, Global Average Pooling, Batch Normalization (BN), and Dropout layers. Furthermore, Optimised training methods including momentum and weight decay were employed. By utilizing deep learning models with varying convolutional kernels during the 1D convolution process, the βGaiaNetβ model effectively captures both local and global features from low-resolution smoothed spectra. Coupled with the nonlinear capabilities provided by an appropriate modelβs depth, the model possesses natural advantages in handling βs XP spectra.
@@
We utilized SHAP for feature attribution interpretation with the βGaiaNetβ classification model based on positive sample. We presented clear feature heatmaps for each spectrum, which serve as a basis for judgment and identification of each spectrum. Based on the statistical analysis of SHAP values from the spectra, we found that the C features provide only a weak signal at the blue end of low flux values and have a minimal presence in many cold carbon stars. Therefore, they were not selected by our model. We summarised five key spectral feature regions: 6780-7460βΓ , 7600-7820βΓ , 7900-8360βΓ , 8820-9060βΓ , and 9140-9560βΓ . Among them, the 7900-8360βΓ Β and 9140-9560βΓ Β regions are the most obvious trough areas. These two regions contribute over 80% of the SHAP value due to CN molecular absorption. The 7600-7820βΓ Β and 8820-9060βΓ Β regions contribute about 16%, which are are also important distinguishing features due to the peak areas from the strong CN molecular absorption on both sides. The 6780-7460βΓ Β region contains relatively weak peaks and troughs, making the smallest contribution less than 4. We hypothesize that this is related to the spectral values, which seems to be addressable through improved spectral normalization methods. These results suggest that our model can effectively learn and capture the key distinguishing features, which are consistent with CN molecular absorption lines and visual identifications by astronomers in the classification of carbon star spectra. These five key feature regions will be essential for future systematic mining of carbon stars from βs XP spectra.
We identified 451 new carbon star candidates from CSTAR_XPM_nonG using the trained classification model. Compared to the band head strength calculation method, our approach effectively detects more carbon star candidates with weaker CN features than those found in the golden sample of carbon stars. These candidates align with the weak CN feature carbon stars depicted in Fig. 21 of C2023. The weaker CN features might be a function of temperature and metallicity effects. Because many of these stars become bluer according to their corrected colors ( Β‘ 2), and show stronger negative contributions caused by CaβII absorption features near 8500βΓ Β based on the statistical analysis of SHAP values compared to the positive sample. Moreover, most of them are not classified as LPVs by Lebzelter etΒ al. (2023), indicating that they could be extrinsic carbon stars. However, the possibility that they are AGB carbon stars with less evolution (limited TPs or mild TDU events) can not be ruled out. Their observed weak carbon spectral feature may suggest a small amount of dredged-up carbon, with C/O ratio not significantly larger than 1. However, caution must be taken when analyzing these candidates, as they are mainly located in the crowded region of Galactic inner disc.
@@
We cross-matched the 451 new candidates with LAMOST DR10, resulting in a total of 4 common sources. Among them, 2 were identified as carbon stars by the LAMOST 1D pipeline. By cross-matching with the SIMBAD database, we obtained 73 common sources. Among these, 28 had already been reported as carbon stars or carbon star candidates in previous work. A reference sample of 1,289 LPVs was constructed to compare the results in Sanders & Matsunaga (2023). In this work, we have discovered 40 potentially carbon stars from it, and they identified 26 C-rich stars including only 6 stars that are not in our candidates. This is due to the significantly different spectral features of the six stars from the positive sample, which results in their confidence lower than 0.5. Furthermore, the assigned confidence and SHAP values for each spectrum were used to quantify the modelβs output, revealing a strong correlation with the band head strengths of R773.3 and R895.0. This suggests that the classification model primarily focuses on the crucial CN features, and the learned features make sense. Hence, our model can efficiently learn the key features between carbon and non-carbon stars in the CSTAR_XPM sample, and the modelβs reliability is further validated.
@@
Due to the limited diversity of the training positive sample, there are still some carbon stars with weaker CN features that are not recognized by our algorithm. Their probabilities given by the modelβs output do not exceed the threshold of 0.5 but significantly larger than 0, thus the proportion of candidates in CSTAR_XPM_nonG could be higher, especially the G and K type stars, predictably if we set a lower threshold. we roughly estimate that the loss percentage is at least about 5%. In future work, it will be feasible to make reasonable improvements to the classification model and incorporate more carbon stars with weaker CN bands into the positive sample to search for and identify more potential warm candidates from CSTAR. We will then expand the algorithm to the entire XP spectral dataset, thereby significantly enriching the number of known carbon stars. Indeed it would be interesting to perform further classification and meaningful analysis. Given the resolution of βs XP spectra, our method can be directly transferred to the upcoming China Space Station Telescope (CSST), especially for the interpretation of different types of stars whose spectral properties are still unclear.
@@ Acknowledgments
Appendix A Heatmaps of six different stars
@@ Appendix B New Carbon star candidates from CSTAR_XPM_nonG
The catalog has 451 entries, which are all the most likely carbon star candidates selected by our model. The model obtains this result by a training positive sample selected from CSTAR_XPM_G and negative sample selected from CSTAR_XPM_non_G. All candidates are sorted in descending order of confidence, and we show the top 100 candidates.